# Achieving more human brain-like vision via human EEG representational alignment Zitong Lu 1,*, Yile Wang 2, Julie D. Golomb1

### v2

@@ 2.4 Improved similarity in ReAlnets to human behavior
Although the results above indicate that ReAlnet, which incorporates additional neural data alignment, demonstrates a higher similarity to human brain representations compared to CORnet, which is purely trained on images, it raises the question: does this neural-level alignment also translate into any behavioral alignment? To test whether ReAlnet shows improved similarity to human behavior, we calculated the scores of CORnet and 10 personalized ReAlnets based on the human behavioral assessments, including two object recognition tasks, in the Brain-Score platform. One task compared how well the ANN, compared to primates, could recognize objects presented in the center of their visual field, even when the objects varied in position, size, viewing angle, and background [7]. The other paradigm tested the similarity of behavioral error between the errors made by humans and ANN on an image-by-image basis [36]. These scores serve as indicators of the models’ similarity to human behavior. The average of two behavioral scores were used to compare with CORnet. Excitingly, the result reveals that ReAlnet, aligned with human EEG data, exhibits representations significantly more akin to human behavior than CORnet does (), further expanding and emphasizing ReAlnet’s status as a more human brain-like vision model.
@@ 2.5 Control experiments
For the control experiments, we tested two aspects: (1) How does contrastive learning influence model-to-brain alignment? (2) If we disrupt the pairing of each image with the EEG signal from the same subject but elicited by viewing a different image, can the model still learn the neural representation patterns of the human brain? Accordingly, we trained two additional sets of ReAlnets based on human EEG data from ten subjects in THINGS EEG2 dataset, termed as W/o ContLoss models (without the constrastive loss component) and Unpaired models (where the pairing between images and EEG signals was disrupted).
@@
We tested the control models on the THINGS EEG2 test dataset and the Shen fMRI dataset, and calculateed the similarity improvement for each control model compared to CORnet. Figure 5 plots the improvement in similarity (compared to CORnet baseline) for ReAlnet and the two controls. Here, we averaged the EEG similarity improvement of all layers and timepoints between 50 and 200ms, and averaged the fMRI similarity improvement of three subjects and five brain regions (See more detailed EEG and fMRI similarity results in Figure S2). The Unpaired control model showed no significant similarity improvement over CORnet for both EEG and fMRI modalities. The W/o ContLoss control model was significantly improved over CORnet for both modalities. W/o ContLoss didn’t perform as well as ReAlnet in terms of fMRI similarity () but showed similar improvement for EEG similarity (). The results of the control experiments reveal (Figure 5A-B): (1) W/o ContLoss models still exhibit an improvement in human brain similarity compared to CORnet. However, while the similarity to human EEG did not decrease compared to ReAlnet, the similarity to cross-modality human fMRI significantly decreased. This suggests that the contrastive loss component in our alignment framework aids ReAlnet in extracting more cross-modality brain visual representation features. (2) Unpaired models failed to enhance brain similarity, which show no significant improvement in brain similarity compared to CORnet, indicating that the training process requires the model to effectively learn the specific neural visual features corresponding to each image. Only in this way can the model become more human brain-like and then exhibit higher similarity to the human brain across different object images, categories, and human neuroimaging data modalities.
@@ 2.6 Human EEG-aligned ResNet also being more brain-like
Although we trained ReAlnet based on CORnet and confirmed that it is more human brain-like, we also wondered whether our multi-layer encoding-based alignment framework could be extended to other models. Therefore, we chose ResNet18, a relatively larger model and aligned it with the EEG representations of ten subjects from THINGS EEG2 dataset using the strategy similar to how we trained ReAlnet. We refer to the aligned model based on pretrained ResNet18 as ReAlnet-R. Subsequentely, we tested ReAlnet-Rs for their similarity to human EEG, fMRI, and behavior, comparing the results with those of the purely image-trained ResNet18.
Firstly, ReAlnet-Rs show significantly higher similarity to human EEG neural dynamics compared to ResNet for nearly all visual layers (Layer 1: 70-160ms, Layer 5: 60-200ms, Layer 9: 60-200ms, Layer 13: 60-180ms, Layer 17: 70-160ms, Figure 6A; See all layers’ EEG similarity results in Figure S3). Secondly, personalized ReAlnet-Rs, similar to ReAlnet, exhibit individual variability increasing with the depth of the layers (Figure 6B; See all layers’ individual variability matrices in Figure S4). Thirdly, ReAlnet-Rs also show higher similarity to human fMRI representations across multiple visual ROIs and different image categories (Figure 6C; See fMRI similarity results on all three subjects in Shen fMRI dataset in Figure S5). Fourthly, for human behavioral similarity, although it is not significant (), seven out of ten ReAlnet-Rs show higher similarity than original ResNet. These results collectively indicate that our alignment framework can be successfully extended to other visual models, such as ResNet, with ReAlnet-R still demonstrating improved similarity to human neural and behavioral representations.
@@ 3 Discussion
Building upon previous research utilizing neural data for aligning object recognition models, we propose a novel and more effective framework for human neural representational alignment, along with the corresponding human brain-like model, ReAlnet. Unlike previous studies that focused on using animal neural signals to optimize models or were unable to use global neural activity for comprehensive model optimization [26, 27, 28, 29, 30], our approach efficiently utilizes human neural activity to simultaneously optimize multiple layers of the model, enabling it to learn the human brain’s internal representational patterns for object visual processing. Notably, unlike prior research relying on behavioral or single modality neural recording data for model evaluation [26, 27, 31, 28, 29, 30], we employed different modalities of human neuroimaging data and also human behaviors for model evaluation to ensure that ReAlnet learns broader, cross-modal brain representational patterns. Additionally, we have extended our alignment framework to another convolutional neural network model to obtain ReAlnet-R and observed a similar enhancement in the similarity to human brain representations.
@@
Regarding ReAlnet itself, it effectively learns not just the patterns of EEG data, but the brain’s internal processing patterns of visual information. This leads to ReAlnet exhibiting a higher similarity than the original CORnet not only to within-modality EEG but also to cross-modality fMRI and behavior. Recent studies in cognitive and computational neuroscience using the THINGS EEG2 dataset have traced evidence of human visual feature processing, including object categories, size, depth, image entropy [37, 38], and even reconstructing visual information through EEG signals [39, 40, 41] and realizing inter-subject EEG conversions [6]. These findings suggest that human brain EEG signals may contain more information than traditional convolutional neural networks can learn from images alone. Similar to this perspective, the fact that different generation loss weights do not significantly impact the model category classification performance but do enhance its similarity to human brains suggests that nodes in the model, which originally did not encode category-specific information, may have been optimized [27]. However, it does warrant further exploration to ascertain what specific information has been learned from the alignment with human brains. More analyses of the neural network’s internal representations may be needed to delve into this. Also, from a reverse-engineering perspective, attempting to understand the brain-like optimization process of the model could further aid in unraveling the mechanisms by which our brains process visual information [42, 43, 44, 45, 46].
Additionally, when we applied this alignment framework to ResNet18, the resulting ReAlnet-R still demonstrates more human brain-like representations, akin to those exhibited by ReAlnet. This framework’s generalizability may suggest that ReAlnet transcends being merely a specific vision model; it represents a pioneering framework potentially applicable for aligning other AI models with brain activity. Therefore, it is worth testing the generalization in the future. On the one hand, this alignment framework may be extended to other neural modalities, such as fMRI and MEG (dimensionality reduction might be necessary for extensive neural data features), paving the way for the development of variants like ReAlnet-fMRI and ReAlnet-MEG. On the other hand, another ambition is to adapt this framework to a wider range of models and tasks in the future, including language and auditory processing and self-supervised or unsupervised models, leading to innovations such as ReAlnet-Language, ReAlnet-Auditory, and self-supervised or unsupervised versions of ReAlnet.
@@
Overall, our study transcends traditional boundaries by employing a groundbreaking alignment framework that pioneers the use of human neural data to achieve a more human brain-like vision model, ReAlnet. Demonstrating significant advances in bio-inspired AI, ReAlnet not only aligns closely with human EEG and fMRI but also exhibits hierarchical individual variabilities and increased similarity to human behavior, mirroring human visual processing. We hope that our alignment framework stands as a testament to the potential synergy between computational neuroscience and machine learning and enables the enhancement of any AI model to be more human brain-like, opening up exciting possibilities for future research in brain-like AI systems.

### v3

@@ 2.4 Improved similarity in ReAlnets to behavior
Does this neural-level alignment also translate into any behavioral alignment? To test whether ReAlnets show improved similarity to human behavior, we calculated the scores of CORnet and 10 personalized ReAlnets based on the human behavioral assessments, including two object recognition tasks, in the Brain-Score platform [37]. One task compared how well the ANN, compared to primates and humans, could recognize objects presented in the center of their visual field, even when the objects varied in position, size, viewing angle, and background [7]. The other paradigm tested the similarity of behavioral error between the errors made by humans and ANN on an image-by-image basis [41]. These scores serve as indicators of the models’ similarity to human behavior. The average of two behavioral scores were used to compare model-behavior similarity. Excitingly, the result reveals that ReAlnets, aligned with human EEG data, exhibit representations significantly more akin to human behavior than CORnet does (ReAlnets vs. CORnet: ; ReAlnets vs. Scrambled models: ) (Figure 4A), further expanding and emphasizing ReAlnets’ status as a more human brain-like vision model. Especially, the results show that ReAlnets exhibit a greater improvement in human behavior similarity on the Geirhos2021 task compared to Rajalingham2018 task. However, the Rajalingham2018 task uses grayscale images where objects are manipulated by varying position, size, viewing angle, and background, while the Geirhos2021 task employs out-of-distribution colorful natural images. A possible explanation is that because ReAlnets are trained with EEG data recorded from subjects viewing natural object images, it better captures human-like behavioral consistency when evaluated on naturalistic, colorful stimuli. Of course, the differences in image manipulation between the two tasks may also contribute to this disparity.
In addition, we submitted our ReAlnets to the updated Brain-Score platform for evaluation and made the scores publicly available on the Brain-Score website (https://www.brain-score.org/vision/). Consistently, the results confirm that ReAlnets achieve significantly higher scores compared to CORnet (see Figure S4), further supporting their improved alignment with brain visual processing.
2.5 Refined object feature representations in ReAlnets
We found that ReAlnets exhibit improved similarity to the human visual system, suggesting that by incorporating human EEG data, they have learned brain representations that purely image-trained models could not capture. Thus, this raises an important question: how do their internal representations differ from those of CORnet? To further investigate what visual object feature representations have been enhanced in ReAlnets by learning human brain signals, we conducted internal representational analysis on purely image-trained CORnet and human EEG-aligned ReAlnets (For the detailed description of the methodology used to analyze these enhanced feature representations, refer to our Methods - Model internal representational analysis section). Here, we tracked feature representations of 49 object feature dimensions from THINGS ([42]) on models’ layer IT to see which features that could be better captured by ReAlnets than CORnet. These 49 object feature dimensions, derived from the THINGS dataset, describe various conceptual and perceptual properties of objects, providing a structured way to assess how different models encode object features beyond simple category labels. Figure 4B shows that ReAlnets show significantly enhanced representations in food-related, artificial, and electronic information (more detailed results of all 49 object feature dimensions are shown in Figure S5). This result suggests that although CORnet trained solely on image data can capture many higher-level object features, some object features, such as food-related information, showed enhanced representation in ReAlnets compared to CORnet. Interestingly, Scrambled models also showed enhanced representations for some of these features, suggesting that even the statistical properties of EEG data can refine internal model representations — though to a lesser extent than real, temporally intact EEG signals.
One might wonder whether these refined object feature representations were simply due to a change in visual ”diet” - that is, the distribution of stimulus categories - when switching from ImageNet to THINGS EEG2 images during alignment. To directly address this potential confound, we conducted a control experiment in which we removed all food-related image-EEG pairs from the training set in THINGS EEG2 prior to the training process. Specifically, we excluded all training images with a ”food” label in THINGS dataset. We then retrained ReAlnets on the reduced dataset using the same alignment framework and evaluated their internal feature representations. Remarkably, even without any exposure to food-related training images or neural signals during our alignment training process, the food-depleted ReAlnets still exhibited significantly enhanced representation of food-related dimensions compared to the original CORnet (Figure S6). While the strength of this enhancement was somewhat reduced, the persistence of food-related representation — despite the complete removal of food stimuli — strongly supports the conclusion that the representational changes observed in ReAlnets are driven primarily by neural alignment rather than by image category frequency.
@@ 2.6 Control experiments
To systematically evaluate how different training manipulations influence model-to-brain alignment, we conducted a set of control experiments. On the one hand, we would like to test the importance of the two loss components in generation loss, reconstruction and contrastive learning losses, in the model alignment. On the other hand, prior non-EEG studies have shown that even optimizing CNNs with shuffled monkey electrophysiology or fMRI signals does not necessarily degrade model performance [26, 43]. Motivated by these, we conducted control experiments to test four aspects: (1) How does the reconstruction loss influence model-to-brain alignment? (2) How does contrastive learning loss influence model-to-brain alignment? (3) If we disrupt the pairing of each image with the EEG signal from the same subject but elicited by viewing a different image, or (4) If we scrambled the EEG time-series data itself, can the model still learn the neural representation patterns of the human brain? Accordingly, we trained four additional sets of ReAlnets based on human EEG data from ten subjects in THINGS EEG2 dataset, termed as W/o ContLoss models (without the constrastive loss component), W/o MSELoss models (without the MSE loss component), Unpaired models (where the pairing between images and EEG signals was disrupted), and Scrambled models (where the EEG time-series were scrambled). More detailed statistical results including t-value, p-value, and Cohen’s d are listed in Tables S2 and S3.
@@
We tested the control models on the THINGS EEG2 test dataset and the Shen fMRI dataset, and calculated the similarity improvement for each control model compared to ReAlnets and the CORnet baseline. Figure 5 plots the improvement in similarity (compared to CORnet baseline) for ReAlnets and the four controls. Here, we averaged the EEG similarity improvement of all layers and timepoints between 50 and 200ms, and averaged the fMRI similarity improvement of three subjects and five brain regions (See more detailed EEG and fMRI similarity results in Figure S7). The W/o MSELoss, Unpaired, and Scrambled control models showed no significant similarity improvement over CORnet for human EEG (W/o MSELoss: ; Unpaired: ; Scrambled: ) and even significant similarity decreasing for human fMRI (W/o MSELoss: ; Unpaired: ; Scrambled: ). The W/o ContLoss control models were significantly improved over CORnet for both modalities (EEG: ; fMRI: ), but they didn’t perform as well as ReAlnets in terms of fMRI similarity () but showed similar improvement for model-EEG similarity ().
The results of the control experiments reveal: (1) W/o ContLoss models still exhibit an improvement in human brain similarity compared to CORnet. However, while the similarity to human EEG did not decrease compared to ReAlnets, the similarity to cross-modality human fMRI significantly decreased. This suggests that the contrastive loss in our alignment framework enables ReAlnets to learn broader and more generalized visual representation patterns across different neuroimaging modalities, such as EEG and fMRI data. (2) The W/o MSELoss, Unpaired, and Scrambled models failed to enhance brain similarity, which show no significant improvement in brain similarity compared to CORnet, indicating that the training process requires the model to effectively learn the specific neural visual features from the actual EEG signals corresponding to each image. Only in this way can the model become more human brain-like and then exhibit higher similarity to the human brain across different object images, categories, and human neuroimaging data modalities.
@@ 2.7 Human EEG-aligned ResNet also being more brain-like
Although we trained ReAlnets based on CORnet and confirmed that they are more human brain-like, we also wondered whether our multi-layer encoding-based alignment framework could be extended to other models. Therefore, we chose ResNet18, a relatively larger model, and aligned it with the EEG representations of ten subjects from THINGS EEG2 dataset using the same framework as above. We refer to the aligned model based on pretrained ResNet18 as ReAlnet-R. Subsequently, we tested ReAlnet-Rs for their similarity to human EEG, fMRI, and behavior, comparing the results with those of the purely image-trained ResNet18.
Firstly, ReAlnet-Rs show significantly higher similarity to human EEG neural dynamics compared to ResNet for nearly all visual layers (Layer 1: 70-160ms, Layer 5: 60-200ms, Layer 9: 60-200ms, Layer 13: 60-180ms, Layer 17: 70-160ms, Figure 6A; Maximum similarity across all layers: 120-180ms, Figure 6B; See all layers’ EEG similarity results in Figure S8). Secondly, personalized ReAlnet-Rs, similar to ReAlnets, exhibit individual variability increasing with the depth of the layers (Figure 6C; See all layers’ individual variability matrices in Figure S9). Thirdly, ReAlnet-Rs also show higher similarity to human fMRI representations across multiple visual ROIs and different image categories (Figure 6D; See fMRI similarity results on all three subjects in Shen fMRI dataset in Figure S10 and detailed statistical results including t-value, p-value, and Cohen’s d are listed in Table S4.). Fourthly, for human behavioral similarity, although it is not significant on the average score (), seven out of ten ReAlnet-Rs show higher similarity than original ResNet. And ReAlnet-Rs show significantly higher behavioral similarity on Geirhos2021 task. These results collectively indicate that our alignment framework can be successfully extended to other visual models, such as ResNet, with ReAlnet-Rs still demonstrating improved similarity to human neural and behavioral representations.
2.8 ReAlnets trained across subjects exhibit higher similarity to EEG and fMRI but not behavior
While the core focus of our study is on individualized ReAlnets that capture subject-specific neural representational patterns, we also evaluated the generalization capability of our alignment framework beyond individual tuning. To this end, we trained an additional model, ReAlnet-AcrossSub, using EEG data pooled across all ten participants in the THINGS EEG2 training set. Concretely, we pooled the EEG data across subjects by treating all trials as coming from a single ”super-subject”, without averaging across them, and the model was trained without subject identifiers. Thus, instead of 10 individualized ReAlnets, we have a single ReAlnet-AcrossSub. Analogous to the single CORnet, we then conducted RSA by comparing the across-subject model against individual subjects’ EEG and fMRI RDMs, as in our main analyses.
ReAlnet-AcrossSub shows significantly higher similarity to human EEG and fMRI representations when human participants viewed natural images compared to CORnet (Figure S11A-C). These results demonstrate that this across-subject model learned unified neural representations across individuals and generalized across neural modalities. However, ReAlnet-AcrossSub did not show model-fMRI similarity improvement on artificial shape or alphabetical letter images. And they did not improve behavioral similarity, performing slightly worse than CORnet (Figure S11D). This dissociation implies that individual-specific neural tuning may be important for improving more brain-like representations of shape or letter visual inputs and capturing human behavioral patterns. We additionally compared the across-subject ReAlnet with the individualized ReAlnets trained separately for each subject (Figure S12). Notably, the relative performance of model-EEG and model-fMRI similarity between them varied slightly across conditions, with individualized ReAlnets performing better in some cases (model-EEG similarity, model-fMRI similarity of shape and letter images, and model-behavior similarity) and the across-subject ReAlnet performing better in others (model-fMRI similarity of natural images).
@@ 3 Discussion
Building upon previous research utilizing neural data for aligning object recognition models, we propose a novel framework for human neural representational alignment, along with the corresponding human brain-like model, ReAlnet. Unlike previous studies that focused on using animal neural signals to optimize models or were unable to use global neural activity for comprehensive model optimization [25, 26, 27, 28, 29], our approach leverages human EEG activity to simultaneously optimize multiple layers of the model, enabling it to learn the human brain’s internal representational patterns for object visual processing. Notably, unlike prior research relying on behavioral or single modality neural recording data for model evaluation [25, 26, 33, 27, 28, 29], we employed different modalities of human neuroimaging data and also human behaviors for model evaluation to ensure that ReAlnets learn broader, cross-modal brain representational patterns. Additionally, we have extended our alignment framework to another convolutional neural network model to obtain ReAlnet-Rs and observed a similar enhancement in the similarity to human brain representations.
Recent advances in brain-inspired AI have explored various approaches to enhancing biological alignment, including self-supervised learning, recurrent architectures, and direct optimization of similarity to neural data. Unlike prior studies that primarily focus on invasive recordings from non-human primates, ReAlnets directly incorporate non-invasive human EEG data, offering a more accessible alternative for modeling human visual representations. While previous work has attempted to improve model-to-brain alignment at a single layer or specific brain region, our multi-layer encoding-based alignment framework enables a more comprehensive alignment across the visual processing hierarchy. This framework is particularly useful in contexts where biological plausibility is a key objective, such as cognitive neuroscience and brain-inspired AI. Recent studies in cognitive neuroscience have begun to apply CNNs to obtain visual features from complex stimuli [44, 45, 46]. Building on this, more brain-like models can now be used to extract representations that are closer to human neural patterns. This approach offers a new way to probe how neural representations envolve across different modalities (e.g., EEG vs. fMRI), stimuli, tasks, and individuals, thereby providing insights into the mechanisms underlying visual perception.
Regarding ReAlnets themselves, they effectively learn not just the patterns of EEG data, but appear to capture something even broader about the brain’s internal processing patterns of visual information. The fact that ReAlnets show higher similarity than the original CORnet not only to within-modality EEG but also to cross-modality fMRI and behavior suggests that the learned representations in ReAlnets capture shared neural patterns that are consistent across different neuroimaging modalities. One possibility is that these shared patterns may reflect the encoding of fundamental visual features, from lower- to higher-level, that are robustly represented in the brain irrespective of the neuroimaging method. In essence, the EEG data could be capturing a core, image-specific representation that generalizes across modalities, providing crucial cues for perceptual inference. And such cross-modal generalization highlights the potential of ReAlnet alignment framework as a flexible framework for extracting the key neural representations beyond single modalities. Our control experiments also highlights that only the model trained on image-specific EEG signals and including both MSE and contrastive learning losses shows this generalization performance. If we remove the contrastive learning loss, we can still see significant model-fMRI similarity improvement but not as great as ReAlnets. If we remove the MSE loss or we employed shuffled (unpaired or time-scrambled) EEG signals to train models, the model-fMRI similarity became even worse than that of the original CORnet. Future work could further explore the specific factors that drive this generalization, such as shared representational structures in EEG and fMRI neural features. And extending this alignment framework to directly incorporate fMRI data or combining fMRI and EEG for joint training could further enhance the model’s ability to capture cross-modal brain representations. Additionally, the ability of ReAlnets to generalize from EEG to fMRI suggests a degree of representational consistency across neural modalities. This observation raises the possibility that different neuroimaging techniques, despite capturing distinct aspects of neural activity (e.g., temporal dynamics in EEG vs. spatial patterns in fMRI), may share a common representational basis [47, 48, 49]. Investigating this consistency further could provide insights into the shared neural mechanisms underlying human visual processing and inform the design of more versatile brain-aligned models.
Recent findings have highlighted the critical role of the training image diet in determining the model-to-brain fit [50]. In addition to the generalization to other neuroimaging modalities, our ReAlnets trained on THINGS EEG2 training dataset also demonstrated robust generalization to held-out THINGS categories. This generalization suggests that the networks can capture shared representational structures that are meaningful both within and beyond the training distribution. Such robustness indicates that training on semantically rich and ecologically valid images and the corresponding brain signals, such as THINGS EEG2, may promote the emergence of brain-like visual representations that generalize across diverse visual domains. On the one hand, our findings demonstrate that our approach improves model-brain similarity. On the other hand, these results may also suggest that selecting appropriate training datasets might be important for letting artificial models learn generalized human brain representations.
@@
EEG signals provide high temporal resolution data that capture rapid neural dynamics underlying visual processing. By leveraging EEG data, our alignment framework enables the model to align with temporal features that likely correspond to distinct stages of visual processing, such as early sensory features and later semantic attributes. Also, compared to fMRI or electrophysiology, EEG is significantly more cost-effective, making it more practical for widespread use. Recent studies in cognitive and computational neuroscience using the THINGS EEG2 dataset have traced evidence of human visual feature processing, including object categories, size, depth, image entropy [38, 51], and even reconstructing visual information through EEG signals [52, 53, 54] and realizing inter-subject EEG conversions [6]. Our internal representational analysis suggests that ReAlnets contain enhanced representations of object dimensions, such as food-related, artificial, electronic, etc. These findings suggest that human brain EEG signals may capture aspects of visual information not present in traditional convolutional neural networks. Similar to this perspective, the fact that different generation loss weights do not significantly impact the model category classification performance but do enhance its similarity to human brains suggests that nodes in the model, which originally did not encode category-specific information, may have been optimized [26]. Also, to address the potential confound that changes in visual “diet” — such as an increased prevalence of food-related items in THINGS compared to ImageNet — may underlie the observed representational shifts, we conducted a control experiment in which all food-related image-EEG pairs were removed during alignment training. Even without exposure to food-related stimuli or neural signals, ReAlnets still exhibited enhanced food-related representations, though to a lesser extent. This suggests that neural alignment, not stimulus statistics alone, is the primary driver of representational changes. While this experiment specifically targeted the food dimension, the findings may reflect a general trend: ReAlnets captures abstract, brain-like representational structure present in neural data, beyond what can be accounted for by training image distributions. However, it does warrant further exploration to ascertain what specific information has been learned from the alignment with human brains. More analyses of the neural network’s internal representations may be needed to delve into this. Also, from a reverse-engineering perspective, attempting to understand the brain-like optimization process of the model could further aid in unraveling the mechanisms by which our brains process visual information [55, 56, 57, 58, 59].
Interestingly, when we applied this alignment framework to ResNet18, the resulting ReAlnet-Rs still demonstrate more human brain-like representations, akin to those exhibited by ReAlnets. This framework’s generalizability may suggest that our alignment framework is potentially applicable to aligning other AI models with human EEG signals. Therefore, one potential direction for future research is to examine whether this alignment framework could generalize to other model architectures or neuroimaging modalities. First, while the current study only utilizes EEG signals for alignment, the observed generalization to fMRI and behavior suggests potential cross-modal consistency in the learned representations. However, we acknowledge that our findings do not directly demonstrate generalization from fMRI to EEG. Future work incorporating direct fMRI-based alignment or joint EEG–fMRI training would be necessary to establish the full bidirectionality of cross-modal generalization. Future studies could test whether this alignment framework could be extended to other neural modalities, such as fMRI and MEG (dimensionality reduction might be necessary for extensive neural data features), suggesting potential applicability of this framework to variants aligned with fMRI or MEG data. Second, another ambition is to adapt this framework to a wider range of models and tasks in the future, including language and auditory processing and self-supervised or unsupervised models, which may be adapted for other modalities and learning paradigms, including auditory or language inputs and unsupervised learning settings. Third, one direction for refinement includes integrating loss functions designed to emphasize task-relevant neural features. Also, although our framework successfully demonstrates generalization to fMRI data, further work is needed to systematically analyze the factors contributing to this transfer. For instance, identifying representational dimensions that are consistent across EEG and fMRI, and relating them to specific brain regions, could provide deeper insights into the mechanisms of cross-modal generalization.
While this study demonstrates encouraging progress, several limitations warrant consideration. First, the relatively small sample size of EEG data and the high level of signal noise may limit the precision of model-to-brain alignment. Second, the lack of shared category labels across datasets—such as the absence of ImageNet labels for THINGS stimuli—complicates consistent model evaluation. These issues may constrain the model’s performance both in brain alignment and object recognition. In addition, although ReAlnets were fine-tuned on EEG signals from individual subjects, they still inherit architectural and representational biases from the pretrained image recognition model, which may limit alignment fidelity. Beyond these data-related factors, certain limitations may also stem from the alignment methodology itself. For instance, our framework employs relatively simple alignment objectives—such as MSE and contrastive losses—that may not fully capture the complex, nonlinear, and distributed nature of brain representations. Furthermore, the shallow encoding modules assume a direct mapping between neural activity and model features, potentially overlooking intermediate transformations or multistage processing. These design choices could limit the model’s capacity to learn more abstract or hierarchically structured neural patterns. Thus, the modest effect sizes observed may reflect not only data constraints, but also the current methodological limits of representational alignment strategies.
Despite these limitations, the framework consistently improves alignment with human EEG, fMRI, and behavior. To build on this work, future efforts could assess whether training the model jointly on EEG and fMRI improves cross-modal generalization, and whether the framework remains robust when using smaller or noisier EEG datasets. Moreover, evaluating how alternative training losses or network architectures impact representational similarity may reveal which components are critical for capturing brain-like representations. Finally, although the THINGS EEG2 dataset is ecologically valid, further work is needed to test whether these findings generalize to datasets with different semantic or contextual structures.
@@
Overall, this study underscores the importance of investigating the limitations and potentials of human EEG signals in shaping brain-like AI. We employ a novel alignment framework using human EEG data to achieve more human brain-like vision models - ReAlnets. Demonstrating significant advances in bio-inspired AI, ReAlnets not only align closely with human EEG and fMRI but also exhibit hierarchical individual variability and increased similarity to human behavior, mirroring human visual processing. We hope that our alignment framework stands as a testament to the potential synergy between computational neuroscience and machine learning and enables the enhancement of diverse AI models to be more human brain-like, opening up exciting possibilities for future research in brain-like AI systems.